OpenStudy (anonymous):

x^2 = 6 + x simplify

OpenStudy (anonymous):

x^2-x-6=0 (x-3)(x+2)=0 x=3,x=-2

OpenStudy (anonymous):

simplify, wasn't sure if I should stop at what I put in here or not

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

no I didn't realise that I had to get two values is all

OpenStudy (anonymous):

many thanks, I'd have been stuck for ages if I'd carried on treating it as a normal equation

OpenStudy (anonymous):

There will always be a number of answers equal to the highest power in your problem. You had an x^2, therefore two answers exist.

OpenStudy (vishweshshrimali5):

i think we just have to simplify it ! so the simplified form i think is x^2-x=6

OpenStudy (vishweshshrimali5):

but if u want me to solve it then i can by quadratic and by factorization @rusty1109

OpenStudy (anonymous):

it's ok, i think I just read the question wrong lol